# CG-HOI: Contact-Guided 3D Human-Object Interaction Generation

### v1

@@ 6.4 Ablation Studies
Figure 8: Application to static scene scans. Our method can generate HOIs from segmented objects in such environments.
6.5 Applications
Human motion generation given object trajectory.
Our approach can also be directly applied to conditionally generate human sequences given object sequences as condition, as shown in Fig. 7. As our model learns a strong correspondence between object and human motion, facilitated by contact distance predictions, we are able to condition without any additional training.
Populating 3D scans.
Fig. 8 shows that we can also apply our method to generate human-object interactions in static scene scans. Here, we use a scene from the ScanNet++ dataset [90], with their existing semantic object segmentation. This enables potential to generate realistic human motion sequences only given a static scene environment.
@@ 6.6 Limitations
While we have demonstrated the usefulness of joint contact prediction in 3D HOI generation, several limitations remain. For instance, our method focuses on realistic interactions with a single object. We show that this can be applied to objects in static 3D scans; however, we do not model multiple objects together, which could have the potential to model more complex long-term human behavior. Additionally, our method requires expensive 3D HOI captures for training; a weakly supervised approach leveraging further supervision from 2D action data might be able to represent more diverse scenarios.
@@ 7 Conclusion
We propose an approach to generating realistic, dynamic human-object interactions based on contact modeling. Our diffusion model effectively learns interdependencies between human, object, and contact through cross-attention along with our contact-based object transformation weighting. Our predicted contacts further facilitate refinement using custom diffusion guidance, generating diverse, realistic interactions based on text descriptions. Since our model learns a strong correlation between human and object sequences, we can use it to conditionally generate human motion sequences from given object sequences in a zero-shot manner. Extensive experimental evaluation confirms both fidelity and diversity of our generated sequences and shows improved performance compared to related state-of-the-art baselines.
@@ Acknowledgements
This project is funded by the Bavarian State Ministry of Science and the Arts and coordinated by the Bavarian Research Institute for Digital Transformation (bidt), and the German Research Foundation (DFG) Grant “Learning How to Interact with Scenes through Part-Based Understanding”.
@@ Appendix B Additional Quantitative Results
B.1 Penetration metric
The exact fidelity and diversity of our results is hard to capture with any single metric. Thus, we evaluate multiple such metrics in the main paper (R-Precision, FID, Diversity, MultiModality), and conduct a perceptual user study to verify the metrics’ expressiveness.
Here, we provide an additional evaluation based on an intuitive physics-based metric: The ratio of frames with human-object inter-penetrations. Due to the imperfect nature of human-object interaction capture, a non-zero amount of penetrations is expected; however, a high amount of penetration indicates low quality interactions with independently floating and often intersecting objects.
We see in Tab. 3 that our approach leads to less overall penetrations, which confirms the higher quality of our sequences, compared to the baselines.
@@ B.1 Penetration metric
InterDiff 0.92 0.81
MDM 0.67 0.78
Ours 0.31 0.36
Table 3: Ratio of frames with penetrations between human and object. Combined with the metrics in the main paper, a lower number corresponds to more realistic interactions. Our approach produces less such inter-penetrations, compared to baselines.
B.2 Novelty of Generated Interactions
@@
Figure 11: Network architecture specification.
@@ F.2 Object Geometry Representation
We represent object geometry as a point cloud, to be processed by a PointNet [56] encoder. For this, we sample $N=256$ points uniformly at random on the surface of an object mesh. Each object category is sampled once as a pre-processing step and kept same for training and inference.

### v2

@@ 6.5 Applications
Populating 3D scans. Fig. 8 shows that we can also apply our method to generate human-object interactions in static scene scans. Here, we use a scene from the ScanNet++ dataset [94], with their existing semantic object segmentation. This enables the potential to generate realistic human motion sequences only given a static scene environment.
Figure 8: Application to static 3D scene scans. Our method can generate HOIs from segmented objects in such environments.
@@ 6.6 Limitations
While we have demonstrated the usefulness of joint contact prediction in 3D HOI generation, several limitations remain. For instance, our method focuses on realistic interactions with a single object. We show that this can be applied to objects in static 3D scans; however, we do not model multiple objects together, which could have the potential to model more complex long-term human behavior (e.g. cooking sequences). Additionally, our method requires expensive 3D HOI captures for training; a weakly supervised approach leveraging further supervision from 2D action data might be able to represent more diverse scenarios. Similarly, our method depends on manual text annotations; more specific prompts might lead to more control over generated HOIs.
@@ 7 Conclusion
We propose an approach to generating realistic, dynamic human-object interactions based on contact modeling. Our diffusion model effectively learns interdependencies between human, object, and contact through cross-attention along with our contact-based object transformation weighting. Our predicted contacts further facilitate refinement using custom diffusion guidance, generating diverse, realistic interactions based on text descriptions. Since our model learns a strong correlation between human and object sequences, we can use it to conditionally generate human motion from given object sequences. Extensive experimental evaluation confirms both fidelity and diversity of our generated sequences and shows improved performance compared to baselines.
@@ Acknowledgements
This project is funded by the ERC Starting Grant SpatialSem (101076253), the Bavarian State Ministry of Science and the Arts coordinated by the Bavarian Research Institute for Digital Transformation (bidt), and the German Research Foundation (DFG) Grant “Learning How to Interact with Scenes through Part-Based Understanding”.
@@ A.1 Additional Interactions
A.2 Same Prompt, Different Interactions
We evaluate the ability of our method to generate diverse interactions for a fixed text condition visually in Fig. 9, for text prompt “Move a stool” and “Sit on a stool”. In the ground truth training data, move is only done with one or two hands, and feet; moving with the butt sometimes occurs for the text description “Sit on a stool”.
Figure 9: Our method is able to generate diverse human-object interactions for the same prompts.
@@ Appendix B Additional Quantitative Results
B.1 Evaluating Penetrations and Floating
Our method discourages penetration and floating implicitly, by enforcing correct contact distances as a soft constraint at train and test time. However, the exact fidelity and diversity of our results is hard to capture with any single metric. Thus, we evaluate multiple such metrics in the main paper (R-Precision, FID, Diversity, MultiModality), and conduct a perceptual user study to verify the metrics’ expressiveness.
Here, we provide an additional evaluation based on intuitive physics-based metrics: Tab. 3 evaluates the mean ratio of frames with some penetration as well as the ratio of penetrating vertices overall, showing that penetrations typically happens with small body parts (e.g., hands, which also occurs in the ground-truth data). We also evaluate the ratio of frames and vertices with human and object not in contact, including floating and stationary objects, which is expected to be close to the dataset ratio.
Results show similar penetration and floating between our generations and ground-truth training data.
@@ B.1 Evaluating Penetrations and Floating
Penetration Non-Contact Penetration Non-Contact
Frames Vertices Frames Vertices Frames Vertices Frames Vertices
Dataset 32.9% 4.1% 21.4% 86.2% 26.9% 1.1% 11.9% 70.4%
Ours 31.3% 3.0% 17.8% 93.3% 35.8% 4.2% 14.1% 74.3%
Table 3: Penetration and non-contact (including floating) ratios in terms of frames as well as overall vertices vs ground-truth data.
B.2 Evaluating Contact
Tab. 4 evaluates our contact predictions using precision/recall and distance metrics. We follow [29, 100, 104] to define contact if ${\leq}5$cm from object. We also report mean $\ell_{1}$ error in contact distance predictions. All metrics are reported for body parts ${\leq}1m$ of the object, to focus on contact scenarios. Better contact prediction corresponds with better HOI generations. Note that none of our baselines predict contact distances.
BEHAVE CHAIRS
Approach Precision $\uparrow$ Recall $\uparrow$ Distance $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Distance $\downarrow$
Separate contact pred. 23.4% 25.6% 0.53 58.6% 49.1% 0.24
No contact weighting 29.5% 33.5% 0.34 60.6% 63.4% 0.10
No contact guidance 46.3% 39.2% 0.31 64.2% 70.2% 0.12
Ours 63.6% 59.5% 0.07 78.3% 84.5% 0.04
Table 4: Evaluation of predicted contact distances, in terms on precision and recall ($\leq 5cm$ distance), as well as mean contact $\ell_{1}$ error in meters.
B.3 Novelty of Generated Interactions
@@
B.4 SMPL Bodies vs. HumanML3D Skeletons
We observe slight pose jitter and foot skating in our ground-truth training data (especially BEHAVE, captured with Kinect sensors). As a result, our model reflects some of these effects. Skeleton representations such as HumanML3D [26] could tackle these artifacts, but do not work with contact as effectively as SMPL bodies. Nevertheless, we train ours with HumanML3D parameters for comparison in Tab. 5 (fitting SMPL after for comparable evaluation) which leads to degraded performance due to less effective contact guidance.
BEHAVE CHAIRS
Representation R-Prec. (top-3) $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ MModality $\rightarrow$ R-Prec. (top-3) $\uparrow$ FID $\downarrow$ Diversity $\rightarrow$ MModality $\rightarrow$
Ours (HumanML3D) 0.33 11.94 2.15 3.75 0.48 12.83 4.39 5.11
Ours 0.62 6.31 6.63 5.47 0.74 6.45 8.91 5.94
Table 5: Ours (using SMPL bodies) vs. using HumanML3D [26] skeletons and fitting SMPL bodies afterwards. While HumanML3D is designed to reduce jitter and foot skating, it leads to degraded performance in our scenario due to less effective contact guidance.
@@ F.2 Object Geometry Representation
We represent object geometry as a point cloud, to be processed by a PointNet [58] encoder. For this, we sample $N=256$ points uniformly at random on the surface of an object mesh. Each object category is sampled once as a pre-processing step and kept same for train and inference.